Wess is 18 years older than soron. Wess is 2.5 times as old as Soron. how old is each?
Answer by Theo(13342)

w = age of weiss
s = age of soron

w = s + 18
w = 2.5 * s

since they are both equal to w, then they are both equal to each other, so you get:

s + 18 = 2.5*s

subtract s from both sides of the equation to get:

18 = 2.5*s - s = 1.5*s

divide both sides of the equation by 1.5 to get

18/1.5 = s

solve for s to get:

s = 12

soron is 12 years old

weiss is 12 + 18 = 30 years old.

weiss is also 2.5 * 12 = 30 years old.

your answer is that soron is 12 years old and weiss is 30 years old.